OpenStudy (anonymous):

Simplify. The result should involve exactly 1 ratio.

OpenStudy (anonymous):

(1/x - x/ x+1) / (2/x+1 - x-1/x)

OpenStudy (anonymous):

\[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b }\times \frac{ d }{ c }\]

OpenStudy (anonymous):

can you please help?

OpenStudy (anonymous):

simplify it? because I am confused...

OpenStudy (anonymous):

Combine the fractions in the top of the ratio first... \[\frac{ 1 }{ x }-\frac{ x }{ x+1 }\] The LCD is x(x+1) so \[\frac{ 1 }{ x }\left( \frac{ x+1 }{ x+1 } \right)-\frac{ x }{ x+1 }\left( \frac{ x }{ x} \right)\] Now multiply those out and simplify .... then we'll go from there.

OpenStudy (anonymous):

x+1/x / xˆ2+x - xˆ2 / xˆ2 + x. is this correct?

OpenStudy (anonymous):

Yes but I would leave the denominator unmultiplied for later. use...\[\frac{ x+1 }{ x \left( x+1 \right)}-\frac{ x^2 }{ x \left( x+1 \right) }\] but now combine the two fractions since they have common den.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

x+1 - xˆ2 / x(x+1) right?

OpenStudy (anonymous):

Well yes ... I thin you mean (x+1-x^2)/[x(x+1)]. I know this seems picky but the order of operations are a pretty picky on this...it is equivalent to:\[\frac{ x+1-x^2 }{ x \left( x+1 \right) }\]

OpenStudy (anonymous):

Now for sake of convention I would suggest writing the terms of the numerator in descending order of power...so that would be:\[\frac{ -x^2+x+1 }{ x \left( x+1 \right) }\]

OpenStudy (anonymous):

yes, I understand this

OpenStudy (anonymous):

Now follow the same procedure for the fractions in the bottom of the ratio (keeping the denominator factored instead of multiplied out)

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

2xˆ2 + 2 - xˆ2 -1 / x(x+1). Is this correct?

OpenStudy (tkhunny):

Please not that this is written very poorly. (1/x - x/ x+1) / (2/x+1 - x-1/x) I believe you mean: (1/x - x/ (x+1)) / (2/(x+1) - (x-1)/x) That's a whole different ballpark.

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

I agree is the bottom of the ratio:\[\frac{ 2 }{ x+1 }-\frac{ x-1 }{ x }\]

OpenStudy (anonymous):

It is not quite what you wrote at the beginning but it makes the most sense so I am assuming you forgot some parentheses.

OpenStudy (anonymous):

@Paynesdad how do I continue for the down ratio? can you please help?

OpenStudy (anonymous):

Is the bottom ratio the one I listed a couple of lines above? (I can't tell from your post as it seems to be missing some key parentheses.)

OpenStudy (anonymous):

yes, that is the bottom ratio

OpenStudy (anonymous):

Ok so what is the LCD of the bottom of the ratio then?

OpenStudy (anonymous):

x+1?

OpenStudy (anonymous):

Not quite x+1 is part of it but not the whole thing. Look at what you did for the top of the ratio.

OpenStudy (anonymous):

x(x+1) ?

OpenStudy (anonymous):

There you go so what do you need to multiply the top and bottom of the first fraction by so that it has the lcd?

OpenStudy (anonymous):

x/x?

OpenStudy (anonymous):

for the first fraction

OpenStudy (anonymous):

yep and the second?

OpenStudy (anonymous):

x+1?

OpenStudy (anonymous):

well (x+1)/(x+1) yes...so go ahead and multiply those out (not the denominators though) and combine the fractions and tell me what you get.

OpenStudy (anonymous):

2x/xˆ2+x - xˆ2+x/xˆ2 +2. is it correct?

OpenStudy (anonymous):

No the first one is good.

OpenStudy (anonymous):

Now you have\[-\frac{ \left( x-1 \right)\left( x+1 \right) }{ x \left( x+1 \right)}\] ... also don't worry about multiplying out the denominator ...it is best to just leave it "unmultiplied" (or in factored form). What does the numerator multiply to?

OpenStudy (anonymous):

this is the second part of the second fraction right? because I am confused

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok I understand now. it the numerator 2x - (x-1)(x+1)?

OpenStudy (anonymous):

yep there you go

OpenStudy (anonymous):

so multiply that and simplify. what do you get?

OpenStudy (anonymous):

i dont know:(

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

Well multiply (x-1)(x+1)... can you do that?

OpenStudy (anonymous):

xˆ2-1

OpenStudy (anonymous):

@Paynesdad can you please help?

OpenStudy (anonymous):

yes so now you can combine the numerators of the two fractions since their denominators are the same... right?

OpenStudy (anonymous):

what do you mean by combine? add them or ...?@Paynesdad

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

yes add, subtract them do what every the expression indicates to do.

OpenStudy (anonymous):

-xˆ2+x-1 + 2x - xˆ2 -1 / x(x+1). is this correct?

OpenStudy (anonymous):

whoa too much....let's back up...you had...

OpenStudy (anonymous):

\[\frac{ 2x }{ x \left( x+1 \right) }-\frac{ x^2-1 }{ x \left( x+1 \right) }\] after multiplying out each fraction right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so when you combine those fractions you subtract the numerators and the denominator stays the same...just like fractions with numbers...so what is that

OpenStudy (anonymous):

2x-xˆ2-1/ x(x+1)

OpenStudy (anonymous):

good now again I would suggest putting the numerator's terms in descending power order. So

OpenStudy (anonymous):

Well wait one thing ... it was 2x - (x^2-1), right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well this why those parentheses are so important...2x-(x^2-1) does not equal 2x - x^2 -1 does it?

OpenStudy (anonymous):

no, it equals 2xˆ4 -2 right?

OpenStudy (anonymous):

No that s way off... remember to distribute the minus sign on the stuff in parentheses!...With \[\frac{ 2x }{ x \left( x+1 \right) }-\frac{ x^2-1 }{ x \left( x+1 \right) }\] you are not just subtracting the x^2 but x^2 and the -1! so\[\frac{ 2x }{ x \left( x+1 \right) }-\frac{ x^2-1 }{ x \left( x+1 \right) }=\frac{ 2x-\left( x^2-1 \right) }{ x \left( x+1 \right) }\] so \[\frac{ 2x }{ x \left( x+1 \right) }-\frac{ x^2-1 }{ x \left( x+1 \right) }\neq\frac{ 2x- x^2-1 }{ x \left( x+1 \right) }\] do you see what I am getting at?

OpenStudy (anonymous):

@RH ?

OpenStudy (anonymous):

is the answer (x-1)ˆ2 / (x+2)ˆ2

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

No...you seem to have gone off the tracks...let's go back to\[\frac{ 2x }{ x\left( x+1 \right) }-\frac{ x^2 -1}{ x \left( x+1 \right) }=\frac{ 2x- \left( x^2-1 \right)}{x \left( x+1 \right) }\] when you distribute the "-" sign what do you get?

OpenStudy (anonymous):

-2xˆ4 + 2x? @Paynesdad

OpenStudy (anonymous):

How are you getting \[-2x^4\] ???

OpenStudy (anonymous):

-2xˆ3 sorry. is it correct?

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

No because you are not multiplying! That is subtraction so you distribute the minus sign to get\[\frac{ 2x -\left( x^2-1 \right)}{ x \left( x+1 \right)}=\frac{ 2x-x^2+1 }{ x \left( x+1 \right) }\]

OpenStudy (anonymous):

Not sure why you were multiplying.

OpenStudy (anonymous):

ok now I understand, sorry for my mistake. how do i continue to simplify?

OpenStudy (anonymous):

:-) ok I would put the numerator in power order just for the sake of convention...so that would be\[\frac{ -x^2 +2x+1}{ x \left( x+1 \right) }\] so now lets put these back into the ratio...So we have...

OpenStudy (anonymous):

-xˆ2 + x - 1/ x(x+1) divided by -xˆ2 + 2x -1/ x(x+1)

OpenStudy (anonymous):

Yes! so now reciprocate the fraction in the bottom of the ratio and cancel what can be and what are you left with?

OpenStudy (anonymous):

I am not sure I understand what you mean with that? :(

OpenStudy (anonymous):

Ok first of all I think you have a couple of -1 that should be +1 double check that

OpenStudy (anonymous):

where???? @Paynesdad

OpenStudy (anonymous):

So you should have \[\frac{ \frac{ -x^2+x+1 }{ x \left( x+1 \right) } }{ \frac{ -x^2+2x+1 }{ x \left( x+1 \right) } }\] ... look closely at everything and make sure you agree

OpenStudy (anonymous):

@RH

OpenStudy (anonymous):

yes!!!! you are right!!! the +1! sorry

OpenStudy (anonymous):

Ok so you can reciprocate the bottom of the ratio and get...

OpenStudy (anonymous):

\[\left( \frac{ -x^2+x+1 }{ x \left( x+1 \right) } \right)\left( \frac{ x \left( x+1 \right) }{ -x^2+2x+1 } \right)\]

OpenStudy (anonymous):

do i cancel out now the same x(x+1)?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

and how do I simplify after that?

OpenStudy (anonymous):

Unless you can factor the polynomials in the numerator and denominator (which you can't) and find common factors to cancel. You are done.

OpenStudy (anonymous):

I would suggest learning the proper place to put parentheses when using / instead of a fraction bar. ... i.e. \[\frac{ 3+2 }{ 4+1 }=(3+2)/(4+1)\] not 3+2/4+1 because \[3+2/4+1=3+\frac{ 1 }{ 2 }+1=4\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

thanks a lot!!!